OpenStudy (anonymous):

A certain merry go round has a radius of 5 meters and an angular velocity of 1/4 pi radians per second. A) how long does it take the merry go round to make one complete revolution? B) How many turns does it make in one minute? I need the equations as well as an explanation

OpenStudy (jdoe0001):

 so how long do you think it'd take for a complete revolution ?

OpenStudy (anonymous):

\[V=r \omega \] \[C=2\pi r\] \[V= (5)(\pi/4)\] V= 3.9 m/s V=d/t \[C=2\pi(5) = 10\pi\] \[3.9=\frac{ 10\pi }{ x }\] x=8 seconds

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Is this correct? Is there a simpler way? Maybe using combined equations?

OpenStudy (anonymous):

How about part B

OpenStudy (jdoe0001):

part B is just saying how many revolutions will it do in 60 secs well, we already know it takes 8secs for 1 revolution, so that'd be simple

OpenStudy (anonymous):

ahh I see 7.5 revolutions

OpenStudy (jdoe0001):

yeap